OpenStudy (anonymous):

Will Medal and Fan to the CORRECT PERSON. line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (daniellelovee):

ok point slope form y-y1=m(x-x1) y+4=-5(x-2)

OpenStudy (daniellelovee):

if they had given you the y intercept then it would have been in the standard slope formula y=mx+b but since they give you points you use point slope formula :)

OpenStudy (anonymous):

thanks.

OpenStudy (daniellelovee):

np

OpenStudy (anonymous):

wait what about the other equation theres more then 1 answer @Daniellelovee